Last year, Goran invested his money in two purchases. He purchased a certificate of deposit for $1000 that paid 2% interest per year and purchased $3000 in corporate bonds paying 7%interest per year.
Answer the questions below. Do not do any rounding.
(a) What was the total interest earned at the end of 1 year?
$ __________________
(b) What was the percent interest for his total investment?
% _________________

Answer by stanbon(75887)

Last year, Goran invested his money in two purchases. He purchased a certificate of deposit for $1000 that paid 2% interest per year and purchased $3000 in corporate bonds paying 7%interest per year.
Answer the questions below. Do not do any rounding.
(a) What was the total interest earned at the end of 1 year?
Int = 0.02*1000 + 0.07*1000 + 0.09(1000) = $90
(b) What was the percent interest for his total investment?
Solve 2000x = 90
x = 90/2000 = 9/200 = 0.045 = 4.5%
_________________
Cheers,
Stan H.